# When Bias Helps Learning: Bridging Initial Prejudice and Trainability

### v3

@@ Abstract
Understanding the statistical properties of deep neural networks (DNNs) at initialization is crucial for elucidating both their trainability and the intrinsic architectural biases they encode prior to data exposure. Mean-field (MF) analyses have demonstrated that the parameter distribution in randomly initialized networks dictates whether gradients vanish or explode. Recent work has shown that untrained DNNs exhibit an initial-guessing bias (IGB), in which large regions of the input space are assigned to a single class. In this work, we provide a theoretical proof linking IGB to MF analyses, establishing that a network’s predisposition toward specific classes is intrinsically tied to the conditions for efficient learning. This connection leads to a counterintuitive conclusion: the initialization that optimizes trainability is systematically biased rather than neutral. We validate our theory through experiments across multiple architectures and datasets.
@@ 1 Introduction
In recent years, deep neural networks have achieved remarkable empirical success across diverse domains [AlphaFold, NEURIPS2020_1457c0d6, pmlr-v139-ramesh21a]. However, understanding their properties theoretically, especially regarding their trainability, remains challenging. A central difficulty consists in explaining how the choice of hyper-parameters — such as weights and biases variances — governs the network’s ability to propagate signals and gradients through depth. Improper initialization typically leads to gradient-related issues: vanishing gradients, causing persistent initial conditions and learning stagnation; or exploding gradients, causing instability in the early stages of training.
A mean-field (MF) theory of wide networks has provided a systematic framework to analyze how these initial parameters shape trainability [schoenholz2016deep, gilboa2019dynamicalisometrymeanfield, pmlr-v80-chen18i, poclópez2024dynamicalmeanfieldtheoryselfattention, yang2019meanfieldtheorybatch, NIPS2017_81c650ca, pmlr-v80-xiao18a, lee2018deepneuralnetworksgaussian, hayou2019impact, jacot2022freeze, poole2016exponential]. Depending on the initialization state, a network exhibits either an ordered phase, where gradients vanish, or a chaotic phase, where gradients explode. The optimal boundary — the so-called "edge of chaos" (EOC) — is characterized by an infinite depth scale in both the forward and backward pass, making the network effectively trainable. This highlights the initial state’s crucial role in determining a network’s subsequent learning dynamics.
Concurrently, fairness has emerged as a central concern [buolamwini2018gender, weidinger2021ethical], driven by the realization that neural networks risk automating discriminatory biases [noble2018algorithms, eubanks2018automating, broussard2018artificial]. Recent insights show that architectural choices significantly impact the behviour of neural networks even before training begins, yielding qualitatively different initial predictive states [pmlr-v235-francazi24a]. Specifically, depending on factors such as network architecture and the initialization of weights, an untrained network may exhibit a prejudice toward certain classes — referred to as initial guessing bias (IGB) — or it may remain neutral, assigning equal frequency to all classes. The impact these initial predictive states have on learnability, however, remains unclear. This begs the question: since IGB is related to initialization, how does it connect to MF theories of initialization?
@@
  • We elucidate the link between predictive initial behaviours (IGB states) and trainability conditions (MF phases), thus connecting learnability and fairness from initialization onward.
  • We show that trainability in wide architectures coincides with a state of transient deep prejudice at initialization, challenging the intuitive assumption [pmlr-v235-francazi24a] that the optimal trainability state must be unbiased.
@@
  • We generalize the IGB framework to accommodate non-zero bias terms and multi-node activation functions (such as maxpool layers), further expanding its applicability, and correct existing MF phase diagram inaccuracies (e.g., for ReLU).
Our results clarify how initialization hyperparameters and architectural choices jointly determine initial predictive behaviours and shape subsequent training dynamics, establishing a clear theoretical connection between initialization, trainability, and fairness considerations.
@@ 2.1 Setup
For the sake of simplicity, we will specialize the discussion to multi-layer perceptrons (MLPs), which still play an important role in modern machine learning as they are the building blocks of most complex architectures. For instance, Transformers [Attention] are composed by several stacked MLPs and attention layers. For MLPs, Eq. (1) reads $\mathcal{W}^{(l)}=\{\boldsymbol{W}^{(l)},\mathbf{b}^{(l)}\}$ and $\mathcal{F}$ is the affine transformation: $\mathcal{F}\bigl(\mathbf{x};\boldsymbol{W}^{(l)},\mathbf{b}^{(l)}\bigr)=\boldsymbol{W}^{(l)}\mathbf{x}+\mathbf{b}^{(l)}$, where $\boldsymbol{W}^{((l))}$ is the weight matrix and $\mathbf{b}^{(l)}$ the bias vector.
@@ 2.2 Order/chaos phase transition: initialization conditions for trainability
In this setup, only one kind of average naturally arises: the average over $\mathcal{W}$ at fixed dataset $\mathcal{D}{}$, which is denoted with an overbar, $\mkern 1.5mu\overline{\mkern-1.5mux\mkern-1.5mu}\mkern 1.5mu\equiv\mathbb{E}_{\mathcal{W}}\left({x|\mathcal{D}}\right)$. For fixed inputs, when performing the limit of infinite width before that of depth, the pre-activations become i.i.d. Gaussian variables with mean $\mu^{(l)}=0$ and signal variance $\sigma^{2}_{y^{(l)}}=q_{aa}^{(l)}$, with $q_{aa}^{(l)}\delta_{ij}=\mkern 1.5mu\overline{\mkern-1.5muy_{i}^{(l)}(a)y_{j}^{(l)}(a)\mkern-1.5mu}\mkern 1.5mu$ [lee2018deepneuralnetworksgaussian]. The infinite-width limit is commonly referred to as mean-field regime, since correlations among neurons vanish and the pre-activation distributions are fully characterized by the signal variance. The pre-activation distributions defined in MF theory permit the study of the propagation of the signal through the network; however, they do not yield any insight into interactions among distinct data samples. To such end, one has to define a correlation coefficient between inputs as $c_{ab}^{(l)}=q_{ab}^{(l)}/\sqrt{q_{aa}^{(l)}q_{bb}^{(l)}}$, where $\mkern 1.5mu\overline{\mkern-1.5muy_{i}^{(l)}(a)y_{j}^{(l)}(b)\mkern-1.5mu}\mkern 1.5mu=q_{ab}^{(l)}\delta_{ij}$, and $q_{ab}^{(l)}$ is the signal covariance between inputs $a$ and $b$. The reader is referred to App. C, where we report the recursive relations for the signal variance (Eq. (7)) and covariance (Eq. (8)), first derived by [poole2016exponential].
@@
The authors of [schoenholz2016deep] extensively analyzed bounded activation functions, such as Tanh. By defining $\chi_{1}\equiv\partial c_{ab}^{(l+1)}/\partial c_{ab}^{(l)}|_{c=1}$, $\chi_{1}=1$ separates an ordered phase ($\chi_{1}<1$) where the correlation coefficient converges to one, i.e. $c\equiv\lim_{l\to\infty}c_{ab}^{(l)}=1$, and a chaotic phase ($\chi_{1}>1$) where the correlation coefficient converges to a lower value. Additionally, the value of $\chi_{1}$ determines the transition from vanishing gradients ($\chi_{1}<1$), to exploding gradients ($\chi_{1}>1$). These two phases have well-known consequences for training: vanishing gradients hinder learning by causing a long persistence of the initial conditions, while exploding gradients lead to instability in the training dynamics [bengio1994learning, pascanu2013difficulty]. At the transition point, both the gradients are stable and the depth-scale of signal propagation diverges exponentially. This is the optimal setting for training, as it allows all layers in the network to be trained from the start.
@@
In this case, the signal variance is not guaranteed to converge, and one has to account for unbounded signals when defining the order/chaos phase transition. For example, it is possible for the correlation coefficient to converge always to one in the whole phase diagram, as we will later demonstrate for ReLU; hence, $c$ does not always serve as an effective order parameter for discriminating between phases. In particular, in App. C we prove that the quantity $\tilde{\chi}_{1}\equiv\partial q_{ab}^{(l+1)}/\partial q_{ab}^{(l)}|_{c=1}$ can discriminate the ordered from the chaotic phase and it is equal to $\chi_{1}$ in the domain of convergence of the variance. Thus, for unbounded activation functions, $\tilde{\chi}_{1}=1$ separates the region in the phase diagram with exploding gradients from the one where gradients vanish, acting as a discriminative order parameter. Following [hayou2019impact], it is therefore appropriate to define the edge of chaos (EOC) as the set of points in the phase diagram where 1) the signal variance converges and 2) $\chi_{1}=\tilde{\chi}_{1}=1$. Additionally, [hayou2019impact] provided a simple algorithm (Algorithm 1 of the main paper) to compute the EOC for a generic single-node activation function. This way, it is also possible to analyze unbounded activation functions in regions of the phase diagram characterized by convergent signals.
@@ 2.3 Initial guessing bias: predictive behaviour at initialization
Within the IGB framework, a predictive bias arises as a consequence of a systematic drift in signal activations. In the presence of IGB, the pre-activation signals at each node are still Gaussian distributed in the infinite-width limit, with variance $\sigma^{2}_{y^{(l)}}$, but each node $i$ is centred around a different point, $\mu_{i}^{(l)}$, which is generally different from zero. The centers only vary with initialization, and are Gaussian-distributed too, with zero mean and variance $\sigma^{2}_{\mu^{(l)}}$. Now the signals related to different nodes in the same layer are distributed differently. This causes a misalignment between the decision boundary — initially positioned near the origin — and the data distribution [pmlr-v235-francazi24a]. Consequently, most input points are assigned to a single class, defining a predictive state which we term as prejudice. Conversely, when the drift is negligible and activations remain symmetrically distributed around zero, predictions remain balanced across classes; we call it neutral state. Prejudice can manifest at different levels, depending on the strength of the classification bias. The extent of activation drift and corresponding predictive bias can be quantified by the activation drift ratio $\gamma^{(l)}$.
@@
Classification fractions
In classification tasks, predictive bias can be quantified by measuring the fraction of inputs, $G_{c}$, classified into each class $c$ at initialization. For illustrative clarity, we consider the binary setting, where the predictive imbalance is fully captured by the fraction of inputs assigned to one reference class. Here, we derive an implicit formula to compute the distribution over $\mathcal{W}$ of the fraction of points classified as a reference class in the infinite-width regime, where all pre-activations are normally distributed:
$G_{0}\equiv\mathbb{P}\left(y_{1}^{(L)}>y_{2}^{(L)}\mid\delta(\mathcal{W})\right)=\Phi\left(\sqrt{\frac{\gamma^{(L)}}{2}}\delta\right),$ (3)
where $\Phi\left(\cdot\right)$ is the Gaussian cumulative function, and $\delta$ is a standard Gaussian variable. The proof is straightforward upon utilizing the IGB pre-activation distributions (see App. D).
The value of $\gamma^{(L)}$ permits to distinguish between three phases and connects them to the prejudice/neutrality framework provided before. When $\gamma^{(L)}\ll 1$, the distribution of $G_{0}$ converges, in the distribution sense, to a Dirac-delta centred in $0.5$, whereas for $\gamma^{(L)}\gg 1$ its distribution converges to a mixture of two Dirac-deltas centred in 0 and $1$, respectively. Remarkably, for $\gamma^{(L)}=1$ (Fig. 1 - middle), $G_{0}$ is uniformly distributed in $(0,1)$; this critical threshold delineates a phase where the fraction of points classified to a reference class exhibits a Gaussian-like shape centred in $0.5$ (Fig. 1 - left), from one where the distribution deviates markedly from Gaussian behaviour and it is bi-modal (Fig. 1 - right). Consequently, neutrality emerges for $\gamma^{(L)}<1$, whereas for $\gamma^{(L)}>1$ the network exhibits prejudice. Moreover, prejudice can compound with depth — we call this deep prejudice — when $\gamma\equiv\lim_{L\to\infty}\gamma^{(L)}=\infty$, resulting in a network manifesting strongly-biased predictions.

### v4

# arXiv:2505.12096v4[cs.LG] 28 Feb 2026

@@ Abstract
The statistical properties of deep neural networks (DNNs) at initialization play an important role to comprehend their trainability and the intrinsic architectural biases they possess before data exposure. Well established mean-field (MF) theories have uncovered that the distribution of parameters of randomly initialized networks strongly influences the behavior of the gradients, dictating whether they explode or vanish. Recent work has showed that untrained DNNs also manifest an initial-guessing bias (IGB), in which large regions of the input space are assigned to a single class. In this work, we provide a theoretical proof that links IGB to previous MF theories for a vast class of DNNs, showing that efficient learning is tightly connected to a network’s prejudice towards a specific class. This connection leads to a counterintuitive conclusion: the initialization that optimizes trainability is systematically biased rather than neutral.
@@ 1 Introduction
In recent years, deep neural networks have achieved remarkable empirical success across diverse domains (AlphaFold; NEURIPS2020_1457c0d6; pmlr-v139-ramesh21a). However, understanding their properties theoretically, especially regarding their trainability, remains challenging. A central difficulty consists in explaining how the choice of hyper-parameters — such as weights and biases variances — governs the network’s ability to propagate signals and gradients through depth. Improper initialization typically leads to gradient-related issues: vanishing gradients, causing persistent initial conditions and learning stagnation; or exploding gradients, causing instability in the early stages of training.
Mean-field (MF) theories of wide networks have provided a systematic framework to analyze how these initial parameters shape trainability (schoenholz2016deep; gilboa2019dynamicalisometrymeanfield; pmlr-v80-chen18i; poclópez2024dynamicalmeanfieldtheoryselfattention; yang2019meanfieldtheorybatch; NIPS2017_81c650ca; pmlr-v80-xiao18a; lee2018deepneuralnetworksgaussian; hayou2019impact; jacot2022freeze; poole2016exponential; pmlr-v139-yang21c). Depending on the initialization state, a network exhibits either an ordered phase, where gradients vanish, or a chaotic phase, where gradients explode. The optimal boundary — the so-called edge of chaos (EOC) — is characterized by an infinite depth scale in both the forward and backward pass, making the network effectively trainable. This highlights the initial state’s crucial role in determining a network’s subsequent learning dynamics.
Beyond their seemingly simplistic theoretical assumption – that is of infinite width – MF theories have led to surprising results valid for real-world models. For the first time, they allowed training of very deep CNN from scratch (pmlr-v80-xiao18a) and hyper-parameters transferability across depth and width through the Tensor Program approach (NEURIPS2019_5e69fda3; yang2020tensorprogramsiineural; pmlr-v139-yang21c). In particular, pmlr-v139-yang21c showed that MF networks are capable of feature learning, effectively exiting the lazy learning regime predicted by the Neural Tangent Kernel (NTK) parametrization. These results underscore the importance of MF theories in effectively describing production architectures, which are in many cases sufficiently large to lay in the MF description.
While previous MF theories specialized on the trainability conditions, recent insights – also derived in the infinite-width limit – showed that architectural choices also impact the initial predictive states (pmlr-v235-francazi24a), significantly changing the classification behaviour of neural networks even before being exposed to data. Specifically, an untrained network may exhibit a prejudice toward certain classes — referred to as initial guessing bias (IGB) — or it may remain neutral, assigning equal frequency to all classes. Surprisingly, IGB solely depends on factors such as network architecture and the initialization of weights. The impact these initial predictive states have on learnability, however, remains unclear. Since IGB is itself a theory of wide-untrained neural networks, it begs the question: how does it connect to previous MF theories of initialization?
Main Contributions
@@
  • We elucidate the link between predictive initial behaviours (IGB states) and trainability conditions (MF phases), clarifying how initialization hyperparameters and architectural choices jointly determine initial predictive behaviours and shape subsequent training dynamics.
  • We show that trainability in wide architectures coincides with a state of transient deep prejudice at initialization, challenging the intuitive assumption (pmlr-v235-francazi24a) that the optimal trainability state must be unbiased.
@@
  • We generalize the IGB framework to accommodate non-zero bias terms and multi-node activation functions (such as maxpool layers), further expanding its applicability, and correct existing MF phase diagram inaccuracies (e.g. for ReLU).
  • We emphasize the role of per-class gradients, showing that the vanishing/exploding gradient behaviour is class-dependent.
Our results are valid for a general wide architecture and only assume its output to be normally distributed, a simpler and more general setting than the original, full IGB analytical framework.
Practical Consequences of Our Theory
Aside from the theoretical value of linking bias with trainability, there are some practical takeaways which become obvious in light of our theory:
  • Gradient exploding. Due to initialization biases, gradient exploding only involves a subset of classes (Fig. 4). This causes an imbalance in per-class gradients, which can drastically slow down learning (francazi2023theoretical).
  • Hyperparameter tuning. Assessing model performances based on short runs is at risk of privileging specific classes due to residual bias (Fig. 5).
  • Knowledge transfer. By linking IGB to previous MF theories, which do not impose any prior data distribution, we relax the assumptions of IGB theory, thus extending its validity. On the converse, the IGB formalism complements the MF phase diagrams, revealing novel phases and unveiling unexpected biases (Fig. 3).
@@ 2.1 Setup
For the sake of simplicity, we will specialize the discussion to multi-layer perceptrons (MLPs), which still play an important role in modern machine learning as they are the building blocks of most complex architectures. For instance, Transformers (Attention) are composed by several stacked MLPs and attention layers. For MLPs, Eq. (1) reads $\mathcal{W}^{(l)}=\{\boldsymbol{W}^{(l)},\mathbf{b}^{(l)}\}$ and $\mathcal{F}$ is the affine transformation: $\mathcal{F}\bigl(\mathbf{x};\boldsymbol{W}^{(l)},\mathbf{b}^{(l)}\bigr)=\boldsymbol{W}^{(l)}\mathbf{x}+\mathbf{b}^{(l)}$, where $\boldsymbol{W}^{((l))}$ is the weight matrix and $\mathbf{b}^{(l)}$ the bias vector.
@@ 2.2 Order/chaos phase transition: initialization conditions for trainability
In this setup, only one kind of average naturally arises: the average over $\mathcal{W}$ at fixed dataset $\mathcal{D}{}$, which is denoted with an overbar, $\mkern 1.5mu\overline{\mkern-1.5mux\mkern-1.5mu}\mkern 1.5mu\equiv\mathbb{E}_{\mathcal{W}}\left({x|\mathcal{D}}\right)$. For fixed inputs, when performing the limit of infinite width before that of depth, the pre-activations become i.i.d. Gaussian variables with mean $\mu^{(l)}=0$ and signal variance $\sigma^{2}_{y^{(l)}}=q_{aa}^{(l)}$, with $q_{aa}^{(l)}\delta_{ij}=\mkern 1.5mu\overline{\mkern-1.5muy_{i}^{(l)}(a)y_{j}^{(l)}(a)\mkern-1.5mu}\mkern 1.5mu$ (lee2018deepneuralnetworksgaussian). The infinite-width limit is commonly referred to as mean-field regime, since correlations among neurons vanish and the pre-activation distributions are fully characterized by the signal variance. The pre-activation distributions defined in MF theory permit the study of the propagation of the signal through the network; however, they do not yield any insight into interactions among distinct data samples. To such end, one has to define a correlation coefficient between inputs as
$c_{ab}^{(l)}=q_{ab}^{(l)}/\sqrt{q_{aa}^{(l)}q_{bb}^{(l)}},$ (2)
where $\mkern 1.5mu\overline{\mkern-1.5muy_{i}^{(l)}(a)y_{j}^{(l)}(b)\mkern-1.5mu}\mkern 1.5mu=q_{ab}^{(l)}\delta_{ij}$, and $q_{ab}^{(l)}$ is the signal covariance between inputs $a$ and $b$. The reader is referred to App. C, where we report the recursive relations for the signal variance [Eq. (7)] and covariance [Eq. (8)], first derived by poole2016exponential.
@@
The authors of schoenholz2016deep extensively analyzed bounded activation functions, such as Tanh. By defining $\chi_{1}\equiv\partial c_{ab}^{(l+1)}/\partial c_{ab}^{(l)}|_{c=1}$, $\chi_{1}=1$ separates an ordered phase ($\chi_{1}<1$) where the correlation coefficient converges to one, i.e. $c\equiv\lim_{l\to\infty}c_{ab}^{(l)}=1$, and a chaotic phase ($\chi_{1}>1$) where the correlation coefficient converges to a lower value. Additionally, the value of $\chi_{1}$ determines the transition from vanishing gradients ($\chi_{1}<1$), to exploding gradients ($\chi_{1}>1$). These two phases have well-known consequences for training: vanishing gradients hinder learning by causing a long persistence of the initial conditions, while exploding gradients lead to instability in the training dynamics (bengio1994learning; pascanu2013difficulty). At the transition point, both the gradients are stable and the depth-scale of signal propagation diverges exponentially. This is the optimal setting for training, as it allows all layers in the network to be trained from the start.
@@
In this case, the signal variance is not guaranteed to converge, and one has to account for unbounded signals when defining the order/chaos phase transition. For example, it is possible for the correlation coefficient to converge always to one in the whole phase diagram, as we will later demonstrate for ReLU; hence, $c$ does not always serve as an effective order parameter for discriminating between phases. In particular, in App. C we prove that the quantity $\tilde{\chi}_{1}\equiv\partial q_{ab}^{(l+1)}/\partial q_{ab}^{(l)}|_{c=1}$ can discriminate the ordered from the chaotic phase and it is equal to $\chi_{1}$ in the domain of convergence of the variance. Thus, for unbounded activation functions, $\tilde{\chi}_{1}=1$ separates the region in the phase diagram with exploding gradients from the one where gradients vanish, acting as a discriminative order parameter. Following hayou2019impact, it is therefore appropriate to define the edge of chaos (EOC) as the set of points in the phase diagram where 1) the signal variance converges and 2) $\chi_{1}=\tilde{\chi}_{1}=1$. Additionally, hayou2019impact provided a simple algorithm (Algorithm 1 of the main paper) to compute the EOC for a generic single-node activation function. This way, it is also possible to analyze unbounded activation functions in regions of the phase diagram characterized by convergent signals.
@@ 2.3 Initial guessing bias: predictive behaviour at initialization
Within the IGB framework, a predictive bias arises as a consequence of a systematic drift in signal activations. In the presence of IGB, the pre-activation signals at each node are still Gaussian distributed in the infinite-width limit, with variance $\sigma^{2}_{y^{(l)}}$, but each node $i$ is centred around a different point, $\mu_{i}^{(l)}$, which is generally different from zero. The centers only vary with initialization, and are Gaussian-distributed too, with zero mean and variance $\sigma^{2}_{\mu^{(l)}}$. Now the signals related to different nodes in the same layer are distributed differently. This causes a misalignment between the decision boundary — initially positioned near the origin — and the data distribution (pmlr-v235-francazi24a). Consequently, most input points are assigned to a single class, defining a predictive state which we term as prejudice, to emphasize that this is a statistically skewed assignment of classes before any data is observed or evaluated.
Conversely, when the drift is negligible and activations remain symmetrically distributed around zero, predictions remain balanced across classes; we call it neutral state. Prejudice can manifest at different levels, depending on the strength of the classification bias. The extent of activation drift and corresponding predictive bias can be quantified by the activation drift ratio $\gamma^{(l)}$.
@@
Classification fractions In classification tasks, predictive bias can be quantified by measuring the fraction of inputs, $G_{c}$, classified into each class $c$ at initialization. For illustrative clarity, we consider the binary setting, where the predictive imbalance is fully captured by the fraction of inputs assigned to one reference class. Here, we derive an implicit formula to compute the distribution over $\mathcal{W}$ of the fraction of points classified as a reference class in the infinite-width regime, where all pre-activations are normally distributed: $G_{0}\equiv\mathbb{P}\left(y_{1}^{(L)}>y_{2}^{(L)}\mid\delta(\mathcal{W})\right)=\Phi\left(\sqrt{\frac{\gamma^{(L)}}{2}}\delta\right),$ where $\Phi\left(\cdot\right)$ is the Gaussian cumulative function, and $\delta$ is a standard Gaussian variable. The proof is straightforward upon utilizing the IGB pre-activation distributions (see App. D).
The value of $\gamma^{(L)}$ permits to distinguish between three phases and connects them to the prejudice/neutrality framework provided before. When $\gamma^{(L)}\ll 1$, the distribution of $G_{0}$ converges, in the distribution sense, to a Dirac-delta centred in $0.5$, whereas for $\gamma^{(L)}\gg 1$ its distribution converges to a mixture of two Dirac-deltas centred in 0 and 11, respectively. Remarkably, for $\gamma^{(L)}=1$ (Fig. 1 - middle), $G_{0}$ is uniformly distributed in $(0,1)$; this critical threshold delineates a phase where the fraction of points classified to a reference class exhibits a Gaussian-like shape centred in $0.5$ (Fig. 1 - left), from one where the distribution deviates markedly from Gaussian behaviour and it is bi-modal (Fig. 1 - right). Consequently, neutrality emerges for $\gamma^{(L)}<1$, whereas for $\gamma^{(L)}>1$ the network exhibits prejudice. Moreover, prejudice can compound with depth — we call this deep prejudice — when $\gamma\equiv\lim_{L\to\infty}\gamma^{(L)}=\infty$, resulting in a network with strongly-biased predictions.
@@
Figure 2: Convergence behaviour of the correlation coefficient of ReLU and Tanh for a single MLP with width equal to 10 000 and depth 100. $\sigma^{2}_{b}=0.1$ and $\sigma^{2}_{w}$ varies uniformly from the ordered phase (blue) to the chaotic phase (red). The transition point is $\sigma^{2}_{w}=2.0$ for ReLU and close to it for Tanh. Scatter points indicate the asymptotic values. The inset plots show the convergence rate for the correlation coefficient to ity asymptotic value $c$, always exponential for Tanh and power law for ReLU in the chaotic phase. Solid lines are computed using the IGB approach [Eq. (6)], while shaded areas represent the 90 $\%$ central confidence interval computed using the MF approach [Eq. (2)].